Solve the system by using the substitution method
x=1/4y-5/4
y=2/5x+14/5
(x,y)=( , )

Answer by checkley77(12569)

x=y/4-5/4
y=2/5x+14/5
y=2/5(y/4-5/4)+14/5
y=2/5(y/4)+2/5(-5/4)+14/5
y=2y/20-10/20+14/5
y=2y/20-10/20+4*14/20
y=(2y-10+56)/20
y=(2y+46)/20
20y=2y+46
20y-2y=46
18y=46
y=46/18
y=23/9 ans.
x=(23/9)*1/4-5/4
x=23/36-5/4
x=23/36-(5*9)/36
x=23/36-45/36
x=-22/36
x=-11/18